OpenStudy (anonymous):

No matter how HARD i try word problems still get me-.- The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1070! If three times the number of one-step equations minus twice the number of two-step equations is equal to 110, how many one-step equations auditioned today?

OpenStudy (anonymous):

okay, we need to write two equations based on the information that is given.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

So take the 2nd sentence and turn it into a equation

OpenStudy (anonymous):

Let E1=1-step equation Let E2= 2-step equation

OpenStudy (anonymous):

E1+E2=1070

OpenStudy (anonymous):

okay sorry i had to do something

OpenStudy (anonymous):

np

OpenStudy (anonymous):

do you understand how i got that equation?

OpenStudy (anonymous):

my apologize ill brb

OpenStudy (anonymous):

and yes i do

OpenStudy (anonymous):

don't worry

OpenStudy (anonymous):

great!

OpenStudy (anonymous):

back

OpenStudy (anonymous):

alright, i want you to say the third sentence out loud and while you do this write down what you are saying on a piece of paper.

OpenStudy (anonymous):

see thats the hard part i am not good at writing down what i read unless its language arts. i had a tutor that would come to my house he would write it on the white board and i would still be confused

OpenStudy (anonymous):

okay, i will write it out

OpenStudy (anonymous):

If three times the number of one-step equations minus twice the number of two-step 3 * E1 - 2* E2 = 110 equations is equal to 110

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Now you have two equations and two unknowns, solve away and then answer your answer is what you got for E1

OpenStudy (anonymous):

what would i need to use to solve it

OpenStudy (anonymous):

either substitution or elimination. these are methods of solutions for systems of inequalities

OpenStudy (anonymous):

wouldnt i be eliminating ?

OpenStudy (anonymous):

Our two equations are: E1+E2=1070 AND 3*E1-2*E2=110

OpenStudy (anonymous):

have you done elimination in systems of equations recently?

OpenStudy (anonymous):

yes but thats my worst one. it get it but i dont get it

OpenStudy (anonymous):

do you get substitution?

OpenStudy (anonymous):

not exactly im on the harder parts right now